a. Mach 10 The daily demand is 60 boats Given the following information Task Fig D, E, F G, H reaponse as a percentage rounded to one decimal place
media%2Fe81%2Fe81f631c-3161-4797-9025-4e
Answer #1

LINE BALANCING

Available time per day= 300 minutes

Daily demand= 60 boats

1.Required cycle time (takt time)

=available production time /( required output or demand ) 300/60=5 minutes per boat

2. Number of workstations (theoretical minimum)

=sum of task times/ required cycle time

=1+1+3+1+2+2+1+2+1=14minutes/5 = 2.8 or 3 station

3. rule for tasks arrangement

  • primary rule - prioritize tasks based on the largest number of following tasks.
  • secondary rule in case of tie- prioritize tasks in the order of the longest task time

Other method-Ranked positional weight

4. Starting with first workstation, assign tasks, one at a time, until the sum of the task times is equal to the workstation cycle time ( or till no other tasks can be assigned due to sequence or time restrictions)

5. Repeat for the remaining workstations till all the tasks have been assigned to a station. We try to pack the earliest workstation first with as much as possible tasks before filling the next ones

6. Line efficiency is the time for making one unit divided by allocated time and is determined by the number of workstations

Line efficiency= Sum of all task times/(Actual number of workstations * required cycle time)

7. Idle time at each station = Required cycle time - sum of task times at each station

8. Idle time for the line per cycle= sum of idle times at individual workstations

Activity

Performance time

number of following tasks

A

1

8

C

3

5

E

2

2

F

2

2

B

1

2

D

1

2

H

2

1

G

1

1

I

1

0

14

Required cycle time=5 minutes

Order of allotment

WS1

Order of allotment

WS2

Order of allotment

WS3

1

A,1

3

E.2

7

H,2

2

C,3

4

F,2

8

G,1

5

B,1

6

D,1

9

I,1

sum of task times each station

5

5

4

Idle time

=5-5=0

=5-5=0

=5-4=1

NOTE

The order of allotment shows how we have ordered the tasks one by one to each station. Note that after F, we allot B to workstation 1 and we try to pack the earliest workstation first with as much as possible tasks before filling the next ones

Line efficiency= Sum of all task times/(Actual number of workstations * required cycle time) = 14/(3*5) =0.93333 or 93.33 %

Idle time for the line per cycle= sum of idle times at individual workstations=0+0+1=1 minute

If 400 minutes are available for production

Required cycle time= 400/60= 6.67 minutes

Theoretical minimum number of workstations = 14 /6.67 = 2.1 or 3 workstations

Theoretical efficiency

= Sum of all task times/(theoretical minimum number of workstations * required cycle time)

=14/(3*6.67)=0.6997 0r 69.97

Note

The Theoretical minimum number of workstations is always rounded up to the next largest integer value

And is not always possible to reach when actually tasks are aligned

A. Answer - Fig 3

Fig 3 is the one that shows 3 works stations and correct precedent relationships

Fig 1- G has B as precedent – wrong

Fig 2- H is in workstation2. It should be in workstation 3

B.93.33%

C.1 minute

D.3 stations

E.69.97%